OpenStudy (anonymous):

Factor completely: −3x2 + 6x − 9

OpenStudy (anonymous):

-3x^2+6x-9

mathslover (mathslover):

\[\large{-3x^2+6x-9}\] do u know about middle term splitting method ?

OpenStudy (anonymous):

No can you explain?

OpenStudy (anonymous):

this cannot be factorise it is irreducible.

OpenStudy (anonymous):

it is either −3(x2 − 2x + 3) or or −1(x2 − 6x + 9) or −3x(x2 − 2x + 3) or −3(x2 + 2x − 3)

OpenStudy (anonymous):

yes you are taking common cannot be factorise in the real no although with iota they can be factorised

OpenStudy (anonymous):

iota?

OpenStudy (anonymous):

=-3x^2+6x-9 you have to break the middle term i.e 6x in such a way that their product will be 27(since -3x^2*-9=27) and sum will be = 6 (from 6x) in this case the middle term will be broken into 3x and 9x becoz 3 n 9 will gave product 27 and sum 6 then place 3 n 9 in place of middle term n solve it =-3x^2+6x-9 =-3x^2+3x+9x-9 take common out them =-3x(x-1)+9(x-1) =(-3x+9) (x-1) this should be the answer =)

OpenStudy (anonymous):

yes it is complex no representation and \[\iota=\sqrt{-1}\]

OpenStudy (anonymous):

@Qibtiya 3x+9x =12x not +6x the middle which it should be

OpenStudy (anonymous):

Since 3,6, and 9 are all multiples of 3, you can factorize 3 from the equation =) If you divide each term by 3, you get: 3(-x^2 + 2x -3) Since you almost always want the highest term to be positive, you can factor the negative sign out, too! So the answer would be -3(x^2 - 2x + 3)

OpenStudy (anonymous):

@sami-21 no for sum we take the middle term as it is the two terms which we select or use must give the sum equal to the middle term i.e 6 hope you understand it =)

OpenStudy (anonymous):

@Qibtiya you wrote =-3x^2+3x+9x-9 if i combine them i get -3x^2+12x-9 which is not the origional polynomial you have added 6x from your own maths rule is if you add spmething subtract it as well also if you multiply the factors u got u do not get the origional question it means something wrong !

OpenStudy (anonymous):

the two terms you selected are 3x and 9x which adds up to 12 x not 6x i hope you got ur typo error !

OpenStudy (anonymous):

yeah thankx sami for the correction we can solve it by taking out the common i.e from whole eq =−3(x2 + 2x − 3) now break the middle term product= -3 and sum =+2 so the two possible values are 3 and -1 =−3(x2 + 2x − 3) =-3[x^2-x+3x-3] =-3[x(x-1)+3(x-1)] =-3[(x+3) (x-1)] this is right answer